OpenStudy (anonymous):

hi, i am doing 6th grade math i need help please explain and please dont just give me the answer i am trying to learn something. Which is the exact area of a circle with radius 14cm

jimthompson5910 (jim_thompson5910):

Area = pi*(radius squared) A = pi*r^2 A = pi*(14)^2 A = pi*14*14 A= pi*196 A = 196pi So the exact area is 196pi square cm

OpenStudy (anonymous):

oh! i see where you got that from so, A= area ?

jimthompson5910 (jim_thompson5910):

yes, A is the area